a triangle (uneven) one side is 3x another is 4x and the other is 5x. The question is find the perimeter, IN FEET if x = 8 INCHES.
How do I figure this out???

Answer by rapaljer(4671)

The perimeter of the triangle = sum of the three sides.
P = 3x + 4x + 5x = 12x

If x= 8 inches, then P= 12x = 12(8 inches) or 96 inches.

To answer in feet, you must divide the previous answer by 12, since there are 12 inches in a foot. This would be 96/12 = 8 feet.

R^2 at SCC